OpenStudy (anonymous):

A student has $1298 in her savings account. She withdraws $95 each week. A formula for calulating the amount of money remaining in her account is A = T - 95w, where T dollars is the original amount and w is the number of weeks she has been withdrawing money. Determine the amount of money remaining in her account after 13 weeks.

TheSmartOne (thesmartone):

What is 95*13 = ? \(\sf\Large 95\times 13 = ?\)

TheSmartOne (thesmartone):

We basically need to find \(\sf\Large A = 1298 - 95(13)\)

OpenStudy (anonymous):

thx

TheSmartOne (thesmartone):

ummm yw .-.

OpenStudy (anonymous):

sorry @TheSmartOne i just figured it out and i was running out of time

TheSmartOne (thesmartone):

ok :P